# A Unified Data Augmentation Framework for Low-Resource Multi-Domain Dialogue Generation

### v1

@@ Abstract
Current state-of-the-art dialogue systems heavily rely on extensive training datasets. However, challenges arise in domains where domain-specific training datasets are insufficient or entirely absent. To tackle this challenge, we propose a novel data Augmentation framework for Multi-Domain Dialogue Generation, referred to as AMD2G. The AMD2G framework consists of a data augmentation process and a two-stage training approach: domain-agnostic training and domain adaptation training. We posit that domain corpora are a blend of domain-agnostic and domain-specific features, with certain representation patterns shared among diverse domains. Domain-agnostic training aims to enable models to learn these common expressive patterns. To construct domain-agnostic dialogue corpora, we employ a de-domaining data processing technique used to remove domain-specific features. By mitigating the effects of domain-specific features, the model trained on the de-domained corpora can effectively learn common expression patterns in different domains. Subsequently, we adapt the learned domain-agnostic features to the target domain through domain adaptation training. We conduct experiments on Chinese dialogue datasets from five different domains and show that AMD2G achieves superior performance compared to both direct training on the target domain corpus and collective training on all five domain corpora. Our work underscores AMD2G as a viable alternative solution for low-resource multi-domain dialogue generation. Code and data associated with our work are available on GitHub repository.
@@ 2.2.2 Dictionary Construction
In addition to keywords extracted by the LLM, we utilize existing terminology banks offered by Chinese input method providers, including QQPinyin111http://cdict.qq.pinyin.cn/v1, Baidu111https://shurufa.baidu.com/dict, and SougouPinyin111https://pinyin.sogou.com/dict/. We retrieve the terms specific to each domain from these sources and merge them into the respective domain’s dictionary. Table 1 presents the statistic overview of the dictionary, including its size, coverage ratio concerning the training set, and the number of replaced tokens. Notably, the domain dictionary’s coverage rate for the training corpus exceeds 95%, with a significant reduction in domain-specific terms.
@@ 3.2 Models
To evaluate the effectiveness and robustness of our proposed method, we apply AMD2G to different types of models, including encoder-decoder and decoder-only structures. For the encoder-decoder structure models, we employ the basic Transformer [42] structure as well as Chinese pre-trained language models such as CPT [40] and the Chinese version of BART111https://huggingface.co/fnlp/bart-base-chinese [10]. The decoder-only structure model used in our experiment is the Chinese version of the pre-trained model GPT-2111https://huggingface.co/uer/gpt2-chinese-cluecorpussmall [33].

### v2

@@ Abstract
Current state-of-the-art dialogue systems heavily rely on extensive training datasets. However, challenges arise in domains where domain-specific training datasets are insufficient or entirely absent. To tackle this challenge, we propose a novel data Augmentation framework for Multi-Domain Dialogue Generation, referred to as AMD2G. The AMD2G framework consists of a data augmentation process and a two-stage training approach: domain-agnostic training and domain adaptation training. We posit that domain corpora are a blend of domain-agnostic and domain-specific features, with certain representation patterns shared among diverse domains. Domain-agnostic training aims to enable models to learn these common expressive patterns. To construct domain-agnostic dialogue corpora, we employ a de-domaining data processing technique used to remove domain-specific features. By mitigating the effects of domain-specific features, the model trained on the de-domained corpora can effectively learn common expression patterns in different domains. Subsequently, we adapt the learned domain-agnostic features to the target domain through domain adaptation training. We conduct experiments on Chinese dialogue datasets from five different domains and show that AMD2G achieves superior performance compared to both direct training on the target domain corpus and collective training on all five domain corpora. Our work underscores AMD2G as a viable alternative solution for low-resource multi-domain dialogue generation. Code and data associated with our work are available on an anonymous GitHub repository***https://github.com/misonsky/Amdg.
@@ 2.2.2 Dictionary Construction
In addition to keywords extracted by the LLM, we utilize existing terminology banks offered by Chinese input method providers, including QQPinyin***http://cdict.qq.pinyin.cn/v1, Baidu***https://shurufa.baidu.com/dict, and SougouPinyin***https://pinyin.sogou.com/dict/. We retrieve the terms specific to each domain from these sources and merge them into the respective domain’s dictionary. Table 1 presents the statistic overview of the dictionary, including its size, coverage ratio concerning the training set, and the number of replaced tokens. Notably, the domain dictionary’s coverage rate for the training corpus exceeds 95%, with a significant reduction in domain-specific terms.
@@ 3.2 Models
To evaluate the effectiveness and robustness of our proposed method, we apply AMD2G to different types of models, including encoder-decoder and decoder-only structures. For the encoder-decoder structure models, we employ the basic Transformer [42] structure as well as Chinese pre-trained language models such as CPT [40] and the Chinese version of BART***https://huggingface.co/fnlp/bart-base-chinese [10]. The decoder-only structure model used in our experiment is the Chinese version of the pre-trained model GPT-2***https://huggingface.co/uer/gpt2-chinese-cluecorpussmall [33].